OpenStudy (symone19):

Use the chart to answer the following question. Pounds:4.536 * 10Negative1 Convert: 90 kilograms to pounds 198.4 1,984 19,840 198,400

Arcadiouse (arcadiouse):

198.416

OpenStudy (symone19):

@Arcadiouse can you tell me steps you did yo get that answers

Arcadiouse (arcadiouse):

1 kg = 2.20462262185 lb

Arcadiouse (arcadiouse):

So 90 x 2.20462262185

Arcadiouse (arcadiouse):

= 198.416

OpenStudy (symone19):

Thank You

Arcadiouse (arcadiouse):

You're more than welcome!

OpenStudy (mathstudent55):

\(1 ~pound = 4.536 \times 10^{-1} ~kg\) \(1~pound = 0.4536~kg\) Divide both sides of the equation above by 0.4536. \(\dfrac{1~pound}{0.4536~kg} = \dfrac{0.4536~kg}{0.4536~kg} = 1\) \(\dfrac{1~pound}{0.4536~kg} = 1\) Now you have a conversion factor that equals one. Start with 90 kg and multiply by the conversion factor. Since the conversion factor is equal to 1, multiplying will not change the amount. It will only change the units. \(90~\cancel{kg} \times \dfrac{1~pound}{0.4536~\cancel{kg}} = 198.413 ~pounds\)